# Implicit Regularization of Infinitesimally-perturbed Gradient Descent Toward Low-dimensional Solutions

### v1

@@ 1.2 Overview of our Contributions
In this paper, our ultimate goal is to address the aforementioned failure mechanism of PGD in the over-parameterized models. At the core of our proposed framework lies a signal-residual decomposition: at each iteration, the current point $\mathbf{x}_{t}$ is decomposed into $\mathbf{x}_{t}=\bm{\mathbf{x}}^{\sharp}_{t}+\bm{\mathbf{x}}^{\perp}_{t}$. Here, $\bm{\mathbf{x}}^{\sharp}_{t}$ represents the signal, defined as the projection of $\mathbf{x}_{t}$ onto the implicit region $\mathcal{M}$. Conversely, $\bm{\mathbf{x}}^{\perp}_{t}:=\mathbf{x}_{t}-\bm{\mathbf{x}}^{\sharp}_{t}$ is the residual. The key insight behind this decomposition is that $\bm{\mathbf{x}}^{\sharp}_{t}$, by its definition, resides within the implicit domain $\mathcal{M}$. Therefore, to establish the convergence to an approximate $\mathcal{M}$-SOSP, it suffices to demonstrate the following: (1) the iterates $\mathbf{x}_{t}$ converge to an approximate SOSP, and (2) the residual norm $\left\lVert\bm{\mathbf{x}}^{\perp}_{t}\right\rVert$ remains small.
Efficient convergence to approximate SOSPs
@@
In this work, we show that large perturbations are unnecessary for PGD to effectively escape strict saddle points. More specifically, we prove that, when GD encounters a strict saddle point, a single perturbation of arbitrarily small radius $\gamma>0$, followed by $T=O(\operatorname{poly}\log(1/\gamma))$ iterations of GD updates, escapes from this strict saddle point. This result is an improvement over the analysis of [20], revealing that the perturbation radius affects the convergence rate only poly-logarithmically, and hence, can be made arbitrarily small at the cost of only poly-logarithmic increase in the number of iterations (Theorem 2). More significantly, this result shows that the impact of an isotropic perturbation—necessary for efficiently escaping strict saddle points—on the residual norm $\left\lVert\bm{\mathbf{x}}^{\perp}_{t}\right\rVert$ can be made arbitrarily small, effectively overcoming the fundamental bottleneck of PGD discussed earlier. We refer to this variant of GD as infinitesimally-perturbed GD (IPGD) (Algorithm 1).
Implicit regularization
@@
Indeed, even if the infinitesimal perturbations of IPGD do not cause a significant increase in the residual norm $\left\lVert\bm{\mathbf{x}}^{\perp}_{t}\right\rVert$, there remains the possibility that the iterates may drift away from the implicit region during the standard GD updates, causing the residual norm $\left\lVert\bm{\mathbf{x}}^{\perp}_{t}\right\rVert$ to grow. To analyze this behavior, we provide a fine-grained characterization of the dynamics of the residual norm, proving that, under reasonable assumptions, the growth or shrinkage of $\left\lVert\bm{\mathbf{x}}^{\perp}_{t}\right\rVert$ is governed by a key quantity, which we call deviation rate (Definition 5). A positive deviation rate drives the shrinkage of the residual norm, while a negative deviation rate drives its growth (Proposition 4).
We combine this result with the saddle-escaping property of IPGD to establish its efficient convergence to an approximate SOSP within the implicit region $\mathcal{M}$ (Theorem 5). In particular, we show that IPGD converges to an $\epsilon$-neighborhood of an $\mathcal{M}$-SOSP within $\tilde{O}(1/\epsilon^{2})$ iterations under three key conditions: (1) the closure of the implicit region $\mathcal{M}$ under GD updates (Assumption 3); (2) the initial point being close to $\mathcal{M}$; and (3) the deviation rate being sufficiently small. We argue that these conditions are satisfied—often provably—in a range of learning problems.
Improved convergence with additional structures
We show that the performance of IPGD can be improved (with a local refinement; see Algorithm 2) if the problem enjoys additional structures within the implicit region $\mathcal{M}$. First, we show that if the function satisfies the strict saddle property within $\mathcal{M}$ (Definition 6), then convergence to an approximate $\mathcal{M}$-SOSP point translates into convergence to a local minimum within $\mathcal{M}$ (Proposition 6). Additionally, if a certain regularity condition (Definition 7) holds within $\mathcal{M}$, the convergence rate can be improved to nearly-linear (Theorem 8). A key distinction of these results compared to the existing work is that, unlike existing work, which assumes these desirable structures hold globally, we only require them to hold within the implicit region $\mathcal{M}$, which is assumed to be low-dimensional. These are significantly milder conditions, as we will discuss.
Applications
@@
We rigorously demonstrate the application of our general framework to a well-known and important learning problem: over-parameterized matrix sensing. This problem serves as an ideal test-bed for our approach for two key reasons. First, as discussed in Section 1.1, general-purpose saddle-escaping methods, such as PGD, are unable to converge to the true solution for this problem. Second, existing tailored analyses of GD for this problem typically involve intricate and lengthy derivations (see, e.g., [6, 5]). In contrast, we demonstrate that our proposed general conditions ensuring the near-linear convergence of IPGD can be established for this problem in a more straightforward manner, significantly simplifying the analysis while achieving near-optimal computational and sample complexities (Theorem 31).
@@
Finally, through additional simulations, we demonstrate that the conditions required for the convergence of IPGD are satisfied across a broad range of over-parameterized problems, including sparse recovery, symmetric and asymmetric matrix completion, and symmetric and asymmetric one-bit matrix completion. Notably, we show that in practice, a perturbation radius of $\gamma=1\times 10^{-15}$—which is close to the round-off error in double precision—is sufficient to ensure the efficient convergence of IPGD. This underscores the insight that, while theoretically distinct, IPGD can be viewed as a practical implementation of GD under inexact arithmetic.
@@ 1.3 Related Work
While most results closely related to our work have already been discussed, there exists a substantial body of literature relevant to various aspects of our findings, which we review below.
Saddle-avoiding algorithms
Although second-order optimization methods—such as cubic regularization [26], trust region methods [27], and their efficient variants [28, 29, 30]—are well-known for their ability to efficiently escape saddle points, their applicability to modern large-scale optimization problems is limited by the computational burden of evaluating the full Hessian, Hessian-vector products, or suitable approximations. However, recent research has demonstrated that even first-order methods can escape saddle points under certain conditions. For instance, GD with random initialization avoids saddle points almost surely [31, 32], but it may do so in exponential time [25]. Moreover, under additional geometric assumptions—such as the invertibility of the Hessian—GD can escape strict saddle points at a linear rate [33, 34]. However, this invertibility condition is rarely satisfied in over-parameterized settings. To address this limitation, numerous GD variants have been developed with provable guarantees for efficient saddle-point escape. As discussed earlier, two such methods—perturbed GD and stochastic GD—leverage random perturbations to escape strict saddle points. In addition, more advanced techniques, such as Nesterov’s accelerated GD [35] and adaptive methods [36], achieve faster escape rates, but may still exhibit significant deviations from the implicit region in the over-parameterized settings.
@@
As mentioned before, local-search algorithms in over-parameterized models often exhibit implicit regularization, converging to low-dimensional solutions (e.g. sparse or low-rank) despite the absence of explicit penalties. For example, training a factorized (over-parameterized) matrix model via alternating or greedy updates typically recovers a low-rank solution: it was conjectured in [37], with supporting evidence, that gradient-based optimization on a full-rank matrix factorization converges to the minimum nuclear norm (i.e., lowest-rank) solution, and it was later shown in [38] that increasing the depth of the factorization amplifies this bias toward low-rank solutions. In tandem, other work has connected this behavior to explicit algorithms: the work [39] proved that the continuous gradient flow with any commuting parametrization is equivalent to continuous mirror descent with a related Legendre function which acts as a regularizer, formalizing the intuition that such local methods perform implicit regularization on a transformed space. Coordinate descent methods likewise impart an implicit sparsity bias. Notably, in classification settings, coordinate-wise descent (analogous to a steepest descent under an $\ell_{1}$ geometry) has been shown to converge to the maximum $\ell_{1}$-margin separator (the sparsest solution), whereas standard gradient descent finds the max-$\ell_{2}$-margin solution [40]. Across these studies, seemingly unconstrained over-parameterized searches end up biased toward low-complexity (low-rank or sparse) solutions, a phenomenon observed empirically and proven in theory in certain applications.
@@
However, prior analyses of implicit regularization have been largely instance-specific, limiting their applicability to broader settings. To the best of our knowledge, this work is the first to introduce a framework for studying the implicit regularization behavior of GD for general over-parameterized models.
Desirable landscape properties
In the literature, several favorable landscape properties have been identified that enable first-order methods to achieve strong convergence guarantees. A central example is the strict saddle property, which ensures that every saddle point possesses at least one direction of sufficiently negative curvature. This property allows common first-order methods—such as PGD [20]—not only to escape saddle points but also to converge to local minima. When combined with the benign landscape property, which guarantees the absence of sub-optimal local minima, this implies that any first-order method capable of escaping saddle points is ensured to converge to a global minimum.
@@
Notably, both the strict saddle and benign landscape properties have been shown to hold in a wide range of problems, including matrix sensing and completion [41, 42], robust PCA [42], complete dictionary learning [43, 44], tensor decomposition [19], phase retrieval [45], and certain classes of neural networks [46, 47, 48]. More broadly, the strict saddle property holds generically in smooth optimization problems [18]. Specifically, for a smooth function $f$ on $\mathbb{R}^{d}$ and almost every perturbation vector $\mathbf{v}\in\mathbb{R}^{d}$, the perturbed function $f_{\mathbf{v}}(\mathbf{x})=f(\mathbf{x})-\langle\mathbf{v},\mathbf{x}\rangle$ satisfies the strict saddle property, as guaranteed by Sard’s theorem. Additionally, many of these problems exhibit structural properties akin to local strong convexity, such as restricted strong convexity or the $(\alpha,\beta)$-regularity condition. These properties are known to accelerate the convergence of GD, often yielding linear rates [49].
@@
However, as we will demonstrate in this paper, assumptions such as the strict saddle property, benign landscape, or local strong convexity and regularity can be overly restrictive in over-parameterized regimes. In such settings, the presence of redundant dimensions can hinder convergence. Instead, we focus on a neighborhood of the implicit solution region, where the adverse effects of over-parameterization diminish and more favorable convergence behavior emerges.
@@ 1.4 Notations
The gradient of a scalar-valued function $f(\mathbf{X})$ with a matrix variable $\mathbf{X}\in\mathbb{R}^{d\times d}$ is an $d\times d$ matrix, whose $(i,j)$-th entry is $[\nabla f(\mathbf{X})]_{i,j}=\frac{\partial f(\mathbf{X})}{\partial X_{ij}}$ for $i,j\in[d]$. Alternatively, we can view the gradient as a linear operator satisfying $\nabla f(\mathbf{X})[\mathbf{Z}]=\langle\nabla f(\mathbf{X}),\mathbf{Z}\rangle=\sum_{i,j}\frac{\partial f(\mathbf{X})}{\partial X_{ij}}Z_{ij}$ for any $\mathbf{Z}\in\mathbb{R}^{d\times d}$. Similarly, the Hessian and the third-order derivative of $f(\mathbf{X})$ can be viewed as multi-linear operators defined by
@@ 2 Main Results
Unlike the classical definitions of gradient- and Hessian-Lipschitz functions, which require smoothness conditions to hold globally, Definitions 1 and 2 impose these conditions only within a $\tau$-neighborhood of the implicit region $\mathcal{M}$. These localized gradient- and Hessian-Lipschitz conditions are particularly advantageous when the smoothness constants scale with the dimensionality of the problem. In such cases, the global smoothness constants scale with the ambient dimension $d$, while their localized counterparts depend only on $\dim(\mathcal{M})=k$. Consequently, when the algorithm operates within the neighborhood of the implicit region, the convergence rate is determined by the local smoothness factors, resulting in a significantly faster rate compared to working over the entire space.
@@
Suppose that $f$ is $(L,\mathcal{M},\tau)$-gradient-Lipschitz and $(\rho,\mathcal{M},\tau)$-Hessian-Lipschitz. Given a tolerance parameter $\epsilon>0$, a point $\mathbf{x}\in\mathcal{N}_{\mathcal{M}}(\tau)$ is called $\bm{\epsilon}$-SOSP if $\left\lVert\nabla f(\mathbf{x})\right\rVert\leq\epsilon$ and $\nabla^{2}f(\mathbf{x})\succeq-\sqrt{\rho\epsilon}\ \mathbf{I}$. Additionally, $\mathbf{x}$ is called $\bm{\epsilon}$$-\bm{\mathcal{M}}$-SOSP if $\mathbf{x}$ is $\epsilon$-SOSP and $\mathbf{x}\in\mathcal{N}_{\mathcal{M}}(\epsilon/L)$ where $\epsilon/L\leq\tau$.
Closely related is the notion of approximate strict saddle points, defined as follows:
Definition 4 (Approximate strict saddle point).
Given a tolerance parameter $\epsilon>0$, a point $\mathbf{x}$ is called $\bm{\epsilon}$-strict saddle if $\left\lVert\nabla f(\mathbf{x})\right\rVert\leq\epsilon$ and $\lambda_{\min}\left(\nabla^{2}f(\mathbf{x})\right)<-\sqrt{\rho\epsilon}$, where $\lambda_{\min}(\cdot)$ denotes the minimum eigenvalue.
Intuitively, we say “$\mathbf{x}$ is an approximate SOSP” or “$\mathbf{x}$ is an approximate strict saddle” if $\mathbf{x}$ is an $\epsilon$-SOSP or $\epsilon$-strict saddle with sufficiently a small parameter $\epsilon>0$. An approximate $\mathcal{M}$-SOSP is defined analogously. Let $\mathcal{X}(\epsilon)$ and $\mathcal{X}_{\mathcal{M}}(\epsilon)$ respectively denote the set of all $\epsilon$-SOSPs and $\epsilon$$-\mathcal{M}$-SOSPs of $f$. Moreover, for simplicity and whenever it is clear from the context, we use $\mathcal{X}$ and $\mathcal{X}_{\mathcal{M}}$ to denote $\mathcal{X}(0)$ and $\mathcal{X}_{\mathcal{M}}(0)$, respectively. Our goal is to recover a point from $\mathcal{X}_{\mathcal{M}}(\epsilon)$ for sufficiently small tolerance $\epsilon>0$. By definition, these points are approximate SOSPs and lie within the proximity of the implicit region $\mathcal{M}$.

### v2

@@ 1.3 Overview of Our Contributions
At the core of our framework is a signal–residual decomposition: each iterate $\mathbf{x}_{t}$ is written as $\mathbf{x}_{t}=\bm{\mathbf{x}}^{\sharp}_{t}+\bm{\mathbf{x}}^{\perp}_{t}$, where $\bm{\mathbf{x}}^{\sharp}_{t}$ is the signal, obtained by projecting $\mathbf{x}_{t}$ onto the implicit region $\mathcal{M}$, and $\bm{\mathbf{x}}^{\perp}_{t}:=\mathbf{x}_{t}-\bm{\mathbf{x}}^{\sharp}_{t}$ is the residual. Since $\bm{\mathbf{x}}^{\sharp}_{t}\in\mathcal{M}$, convergence to an approximate $\mathcal{M}$-SOSP requires showing that (i) $\mathbf{x}_{t}$ approaches an SOSP and (ii) the residual norm $\|\bm{\mathbf{x}}^{\perp}_{t}\|$ remains small.
Escaping saddle points efficiently.
@@
We show that large perturbations are unnecessary for GD to escape SSPs. Specifically, we prove that, when GD encounters an SSP, a single perturbation of arbitrarily small radius $\gamma>0$, followed by $T=O(\operatorname{polylog}(1/\gamma))$ iterations of GD updates, enables the algorithm to escape from this SSP. This result is an improvement over the analysis of [jin2017escape], revealing that the perturbation radius affects the convergence rate only poly-logarithmically, and hence, can be made arbitrarily small at the cost of only poly-logarithmic increase in the number of iterations (Theorem 2). Crucially, such infinitesimal isotropic perturbations preserve a small residual norm, overcoming the main limitation of projected GD. We refer to this variant as Infinitesimally Perturbed Gradient Descent (IPGD) (Algorithm 1).
Implicit regularization.
@@
Even with infinitesimal perturbations, the iterates may drift away from $\mathcal{M}$ during the standard GD updates, causing the residual norm $\left\lVert\bm{\mathbf{x}}^{\perp}_{t}\right\rVert$ to grow. We analyze this behavior by tracking the dynamics of $\|\bm{\mathbf{x}}^{\perp}_{t}\|$ and show that its evolution is governed by a quantity termed the deviation rate (Definition 5). A negative deviation rate contracts the residual, while a positive one causes drift (Proposition 1). Combining this with the saddle-escaping property, we establish that IPGD converges to an $\epsilon$-neighborhood of an $\mathcal{M}$-SOSP within $\tilde{\mathcal{O}}(1/\epsilon^{2})$ iterations, under three conditions: (1) closure of $\mathcal{M}$ under GD updates, (2) initialization near $\mathcal{M}$, and (3) a sufficiently small deviation rate (Theorem 3).
Improved convergence under additional structure.
We show that, when the function satisfies the strict-saddle property within $\mathcal{M}$, convergence to an $\mathcal{M}$-SOSP implies convergence to a local minimum (Proposition 2). Moreover, under a mild regularity condition, the convergence rate of IPGD improves to nearly linear (Theorem 4). Unlike prior analyses that assume global structure, our results only require these properties to hold within the low-dimensional manifold $\mathcal{M}$.
Over-parameterized matrix sensing.
@@
We illustrate our framework on a canonical problem: over-parameterized matrix sensing. This problem serves as an ideal test-bed because, as discussed in Section 1.2, general-purpose saddle-escaping methods like PGD fail to reach the true solution, while existing GD analyses are often intricate and lengthy [stoger2021small, li2018algorithmic]. In contrast, our general conditions yield a concise and unified proof of IPGD’s near-linear convergence, greatly simplifying the analysis (Theorem 5).
Broader applications.
@@
Finally, we verify empirically that the assumptions ensuring IPGD convergence hold across various over-parameterized problems, including matrix completion (symmetric, asymmetric, and one-bit) and sparse recovery. Notably, we show that in practice, a perturbation radius of $\gamma=1\times 10^{-15}$—which is close to the round-off error in double precision—suffices to ensure the efficient convergence of IPGD. This underscores the insight that, while theoretically distinct, IPGD can be viewed as a practical implementation of GD under inexact arithmetic.
@@ 1.4 Related Work
Saddle-avoiding algorithms.
Second-order methods such as cubic regularization [nesterov2006cubic], trust-region algorithms [curtis2017trust], and their efficient variants [carmon2018accelerated, agarwal2017finding, carmon2019gradient] are well known for efficiently escaping SSPs. However, their use in large-scale problems is limited by the high cost of computing or approximating Hessian information. Recent work shows that even first-order methods can escape SSPs under suitable conditions: GD with random initialization avoids them almost surely [lee2019first, panageas2019first]; under stronger geometric assumptions, such as global Hessian invertibility, GD can escape linearly [dixit2023exit, dixit2022boundary], though such assumptions rarely hold in over-parameterized regimes.
To overcome these limitations, several GD variants have been proposed. PGD escapes SSPs using episodic random perturbations, a property shared by its non-periodic version [jin2021nonconvex], stochastic GD with dispersive noise [fang2019sharp], and earlier noisy or normalized variants [ge2015escaping, levy2016power]. Such additive noise appears essential, as plain GD may take exponential time to escape [du2017gradient]. More advanced techniques, including accelerated GD [jin2018accelerated] and adaptive methods [staib2019escaping], achieve faster escape rates but can still deviate significantly from the implicit region in over-parameterized settings.
@@
As discussed earlier, local-search algorithms in over-parameterized models often exhibit implicit regularization, converging to low-dimensional solutions (e.g., sparse or low-rank) even without explicit penalties. For instance, training factorized matrix models via gradient or alternating updates typically yields low-rank solutions: [gunasekar2017implicit] conjectured, and provided evidence, that GD on full-rank matrix factorizations converges to the minimum nuclear-norm solution, while [arora2019implicit] showed that increasing the factorization depth amplifies this low-rank bias. Relatedly, [li2022implicit] proved that continuous gradient flow under any commuting parametrization is equivalent to mirror descent with a Legendre function acting as a regularizer, formalizing implicit regularization in transformed spaces. Coordinate descent methods exhibit similar sparsity biases: in classification, coordinate-wise descent converges to the maximum $\ell_{1}$-margin (sparsest) separator, whereas standard GD yields the maximum $\ell_{2}$-margin solution [soudry2018implicit].
@@
Despite these advances, existing analyses of implicit regularization remain instance-specific. To our knowledge, this work is the first to present a unified framework for studying the implicit regularization behavior of GD in general over-parameterized models.
Desirable landscape properties.
@@
Several landscape conditions have been identified that enable first-order methods to achieve strong convergence guarantees. A central one is the strict saddle property, which ensures every saddle point has a direction of negative curvature, allowing algorithms such as PGD [jin2017escape] to escape SSPs and converge to local minima. Combined with benign nonconvexity—the absence of suboptimal local minima—this guarantees global convergence for any method that avoids saddle points. These properties hold in many problems, including matrix sensing and completion [bhojanapalli2016global, ge2017no], robust PCA [ge2017no], dictionary learning [sun2016complete1, sun2016complete2], tensor decomposition [ge2015escaping], phase retrieval [sun2018geometric], and certain neural networks [brutzkus2017globally, du2017convolutional, soltanolkotabi2018theoretical]. More generally, the strict saddle property holds generically in smooth optimization problems [davis2019active]: for a smooth $f$ on $\mathbb{R}^{d}$ and almost every perturbation $\mathbf{v}$, the function $f_{\mathbf{v}}(\mathbf{x})=f(\mathbf{x})-\langle\mathbf{v},\mathbf{x}\rangle$ satisfies this property. Many such problems also exhibit structural regularities, e.g., restricted strong convexity or $(\alpha,\beta)$-regularity, that further accelerate GD [chi2019nonconvex].
@@
However, these assumptions can be restrictive in over-parameterized regimes, where redundant dimensions may hinder convergence. Instead, we focus on neighborhoods of the implicit solution region, where over-parameterization effects vanish and favorable convergence behavior emerges.
@@ 1.5 Notations
The gradient of a scalar-valued function $f(\mathbf{X})$ with a matrix variable $\mathbf{X}\in\mathbb{R}^{d\times d}$ is a $d\times d$ matrix, whose $(i,j)$-th entry is $\nabla f(\mathbf{X})]_{i,j}=\frac{\partial f(\mathbf{X})}{\partial X_{ij}}$ for $i,j\in[d]$. Alternatively, we can view the gradient as a linear operator satisfying $\nabla f(\mathbf{X})[\mathbf{Z}]=\langle\nabla f(\mathbf{X}),\mathbf{Z}\rangle=\sum_{i,j}\frac{\partial f(\mathbf{X})}{\partial X_{ij}}Z_{ij}$ for any $\mathbf{Z}\in\mathbb{R}^{d\times d}$. Similarly, the Hessian and the third-order derivative of $f(\mathbf{X})$ can be viewed as multi-linear operators defined by
@@ 2 Main Results
Unlike the classical definitions of gradient- and Hessian-Lipschitz functions, which require smoothness conditions to hold globally, Definitions 1 and 2 impose these conditions only within a $\tau$-neighborhood of the implicit region $\mathcal{M}$. These localized gradient- and Hessian-Lipschitz conditions are advantageous when the smoothness constants scale with the dimensionality of the problem. In such cases, the global smoothness constants scale with the ambient dimension $d$, while their localized counterparts depend only on $\dim(\mathcal{M})=\dim(\varphi^{-1}(\mathcal{D}))$ which, depending on $\dim(D)$ and the choice of $\varphi$, can be significantly smaller than $d$.
@@
[…]
Closely related is the notion of approximate strict saddle points (SSPs), defined as follows:
Definition 4 (Approximate SSP).
Given a parameter $\epsilon>0$, a point $\mathbf{x}$ is called $\bm{\epsilon}$-SSP if $\left\lVert\nabla f(\mathbf{x})\right\rVert\leq\epsilon$ and $\lambda_{\min}\left(\nabla^{2}f(\mathbf{x})\right)<-\sqrt{\rho\epsilon}$, where $\lambda_{\min}(\cdot)$ denotes the minimum eigenvalue.
We say $\mathbf{x}$ is an approximate SOSP” or $\mathbf{x}$ is an approximate SSP” if $\mathbf{x}$ is an $\epsilon$-SOSP or $\epsilon$-SSP with a sufficiently small parameter $\epsilon>0$. An approximate $\mathcal{M}$-SOSP is defined analogously. Let $\mathcal{X}(\epsilon)$ and $\mathcal{X}_{\mathcal{M}}(\epsilon)$ respectively denote the set of all $\epsilon$-SOSPs and $\epsilon$-$\mathcal{M}$-SOSPs of $f$. For simplicity and whenever it is clear from the context, we use $\mathcal{X}$ and $\mathcal{X}_{\mathcal{M}}$ to denote $\mathcal{X}(0)$ and $\mathcal{X}_{\mathcal{M}}(0)$, respectively. Our goal is to recover a point from $\mathcal{X}_{\mathcal{M}}(\epsilon)$ for sufficiently small tolerance $\epsilon>0$. By definition, these points are approximate SOSPs and lie within the proximity of the implicit region $\mathcal{M}$.